2. Given F(K,L) = AKa Lb

, show that a and b determine whether the function is contact, increasing 

or decreasing return to scale (show one example in each case).

3. Explain two major approaches to the analysis of consumer behavior.

4. How is consumer equilibrium reached (explain in words) ? Mathematically, show how it is 

reached


1
Expert's answer
2021-03-09T07:35:00-0500

F(K,L) = AK"^{a}"L"^{b}"


Using the properties of homogeneity

Multiply each input by a constant, say, "k" , and factor

F("k" K,"k" L) = "A(kK)^a(kL)^b"

= "Ak^aK^ak^bL^b"

= "k^{a+b}(AK^aL^b)"

= "k^{a+b}(F(K,L))"

For a strict Cobb-Douglas production function, where "a+b=1" , exhibits a constant returns to scale. Where "a+b\u22601" (a generalized production function), if "a+b>1" , the function exhibits an increasing returns to scale and a decreasing returns to scale if "a+b<1"


2) Approaches to consumer behavior analysis include the preference approach and utility approch.


The preference approach makes use of the indifference curve to show how indifferent a consumer is to the consumption of different bundles of commodities given his budget constraints.


The utility approach looks at consumer behaviour from the aspect of ranking. The approach seeks to measure the satisfaction derived from the consumption of certain goods by consumers.


3) Consumers equilibrium is achieved at the point where the budget constraints is at tangency with the indifference curve. This gives the optimum bundle available to a consumer given his preference and constraint. At this point, the ratio of the marginal utilities, "\frac{MUx}{MUy}" must equal the price ratio, "\frac{Px}{Py}"

Mathematically, this is derived by taking the marginal benefits of the goods and equating to the marginal limits of the market.


Assume a utility function "u=\u221a(xy)" subject to the constraint "PxX+PyY=B"

Where u=√(xy) = "√(x^{\frac{1}{2}}y^{\frac{1}{2}})"

U="X^{\frac{1}{2}}Y^{\frac{1}{2}}-\lambda(B-Px-Py)"


Ux= "\frac{1}{2}X^{-\frac{1}{2}}Y^{\frac{1}{2}}-\lambda{Px}" = 0

Uy="\frac{1}{2}X^{\frac{1}{2}}Y^{-\frac{1}{2}}-\lambda{Py}" = 0

U"{\lambda}" = "B-Px-Py" = 0


Where "\frac{1}{2}X^{-\frac{1}{2}}Y^{\frac{1}{2}}" = Ux = MUx and "\frac{1}{2}X^{\frac{1}{2}}Y^{-\frac{1}{2}}" = Uy =MUy

"\lambda=\frac{{{\frac{1}{2}}}X^{-\frac{1}{2}}Y^{\frac{1}{2}}}{Px}" = "\frac{MUx}{Px}"

"\lambda=\frac{{{\frac{1}{2}}}X^{\frac{1}{2}}Y^{-\frac{1}{2}}}{Py}" ="\frac{MUy}{Py}"

Equating "\lambda"

"\frac{MUx}{Px}" = "\frac{MUy}{Py}"

"\frac{MUx}{Px}=\frac{MUy}{Py}"